A simple random sample of 40 items results in a sample mean of 25 . The population standard deviation is a=5 . The 95% confidence interval of the population mean is
1. (25 ; 1,5495)
2. (23,4006 ; 26,5993)
3. (23,4505 ; 26,5495)
4. (23,6679 ; 26,3321)
5. (25,516 ; 27,4591)
1
Expert's answer
2020-11-09T20:18:48-0500

We need to construct the 95% confidence interval for the population mean "\mu."

The following information is provided:


"\bar{x}=25, \sigma=5, n=20"

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as 


"CI=(\bar{x}-\sigma\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+\sigma\times\dfrac{\sigma}{\sqrt{n}})="

"=(25-1.96\times\dfrac{5}{\sqrt{40}}, 25+1.96\times\dfrac{5}{\sqrt{40}})="

"=(23.4505, 26.5495)"

"3.\ \ \ (23.4505, 26.5495)"